# Part I: Nonlinear Codes

### v1

@@ III Direct Constructions for de Bruijn Arrays Codes
Note, that when $t=0$, a $(2^{k},2^{t};n,1)$-DBAC is a perfect factor and when $t=n-k$, a $(2^{k},2^{t};n,1)$-DBAC is a de Bruijn array.
Next, we would like to have a direct construction for $(2^{k},2^{t};n,2)$-DBACs. For $t=1$ such codes do not exists by Lemma 1. Hence, we are intereted in the case where $t>1$. We notice that to construct such codes we can take a PF$(n,k)$, say ${\cal F}$, and guarantee that each two ordered sequences of ${\cal F}$ (not necessarily distinct) will be in all possible shifts one towards the other. This can be done in most cases as will be explained in the proof of the next theorem.
Theorem 6
If there exist a PF$(n,k)$, where $1\leq n-k$, then there exists a $(2^{k},2^{t};n,2)$-DBAC, where $k<n<2^{k}$ and $12\leq t\leq 2n-k-$ 1 (if $t=2n-k$ then we have a $(2^{k},2^{t};n,2)$-DBA).
Proof:
Let ${\cal F}$ be a PF$(n,k)$ which by definition has $2^{n-k}$ cycles. To construct the appropriate code $\mathbb{C}$ of size $2^{2n-t-k}$ we have to show that we can arrange the sequences of ${\cal F}$ as columns in $2^{k}\times 2^{t}$ arrays such that each $n\times 2$ matrix is a contained as an $n\times 2$ window in exactly one of these arrays. Since each $n$-tuple appears as a window of length $n$ in one of the sequences of ${\cal F}$, it follows that each ordered two sequences of ${\cal F}$ (not necessarily distinct) must be in all the possible shifts between them. For each such a pair of sequences $({\cal S}_{1},{\cal S}_{2})$, we must have for each $i$, $10\leq i\leq 2^{k}-$ 1, the ordered pair $({\bf E}^{j}{\cal S}_{1},{\bf E}^{j+i}{\cal S}_{2})$ is contained as a pair of columns in exactly one of the codewords of $\mathbb{C}$ for some $j$. First, we will show that this is possible in the sense that the sum of all the shifts is 0 modulo $2^{k}$. Between any ordered pair of sequences the shifts are from 0 to $12^{k}-$ 1. The number of such ordered pairs is $2^{n-k}\cdot 2^{n-k}$. Hence, the total sum of all the shifts is
$\left(2^{n-k}\cdot 2^{n-k}\right)\sum_{i=0}^{2^{k}-1}i=2^{2(n-k)}(2^{k}-1)2^{k-1}\equiv 0~{}({\mbox{mod}}~{}2^{k})~{},$
since $n-k\geq 1$.
  1. 1.
    Assume first that $k=n-1$, i.e., ${\cal F}$ contains two sequences ${\cal A}$ and ${\cal B}$ of length $2^{n-1}$. We distinguish now between three cases depending on the value of $t$.
    Case 1: If $t=n$ (i.e., $t=k+1$), then there are exactly two codewords in $\mathbb{C}$ and they can be constructed as follows.
    $[{\cal A},{\bf E}{\cal B},{\cal A},{\bf E}^{2}{\cal B},{\cal A},{\bf E}^{3}{\cal B},\ldots,{\cal A},{\bf E}^{2^{t-1}-2}{\cal B},{\cal A},{\bf E}^{2^{t-1}-1}{\cal B},{\bf E}^{2^{t-1}-1}{\cal B},{\cal A}]$ (3)
    $[{\cal A},{\bf E}{\cal A},{\cal A},{\bf E}^{2}{\cal A},\ldots,{\cal A},{\bf E}^{2^{t-2}-1}{\cal A},{\cal A},{\bf E}^{2^{t-2}}{\cal A},{\bf E}^{2^{t-2}}{\cal B},{\cal B},{\bf E}^{2^{t-2}-1}{\cal B},{\cal B},\ldots,{\bf E}^{2}{\cal B},{\cal B},{\bf E}{\cal B},{\cal B}]~{}.$ (4)
    In the first codeword we only have the pairs $({\cal A},{\cal A})$ and $({\cal B},{\cal B})$ (we view the first sequence as being in its zero state) with the same two sequences in the pair. In the second codeword we have all the pairs of the form $({\cal A},{\bf E}^{i}{\cal A})$, where $1\leq i\leq 2^{t-2}$, and $({\bf E}^{i}{\cal A},{\cal A})$ which is equivalent to $({\cal A},{\bf E}^{-i}{\cal A})$, where $11\leq i\leq 2^{t-2}-$ 1. Since $k=t-1$ and the shifts are taken modulo $2^{k}$, it follows that each pair $({\cal A},{\bf E}^{i}{\cal A})$, where $0\leq i<2^{k}$, appears exactly once in one of the two codewords. Exactly the same arguments holds for the pairs of the form $({\cal B},{\bf E}^{i}{\cal B})$.
    In the second codeword we only have the pairs $({\cal A},{\cal B})$ and $({\cal B},{\cal A})$, where the two sequences in the pair are distinct. In the first codeword we have all the pairs of the form $({\cal A},{\bf E}^{i}{\cal B})$ and $({\bf E}^{i}{\cal B},{\cal A})=({\cal B},{\bf E}^{-i}{\cal A})$, where $11\leq i\leq 2^{t-1}-$ 1. Each such pair appears exactly once and since $k=t-1$, it follows that each required pair appears exactly once in one of the two codewords.
    Thus, $\mathbb{C}$ is a $(2^{k},2^{t};n,2)$-DBAC.
    Case 2: If $t=n-1$ (i.e., $t=k$), then there are exactly four codewords in $\mathbb{C}$. The first two codewords can be defined as in Eqs. (3) and (4), but the shifts are smaller since $t$ is smaller compared to $t$ of Case 1. This also implies that the number of columns in the codeword are smaller. The other two codewords can be defined as follows.
    $[{\cal B},{\bf E}^{2^{t-1}-1}{\cal A},{\cal B},{\bf E}^{2^{t-1}-2}{\cal A},{\cal B},{\bf E}^{2^{t-1}-3}{\cal A},\ldots,{\cal B},{\bf E}^{2}{\cal A},{\cal B},{\bf E}{\cal A},{\bf E}^{2^{k-1}+1}{\cal A},{\bf E}^{2^{k-1}}{\cal B}]$ (5)
    $[{\cal B},{\bf E}^{2^{t-1}-1}{\cal B},{\cal B},{\bf E}^{2^{t-1}-2}{\cal B},\ldots,{\cal B},{\bf E}^{2^{t-2}+1}{\cal B},{\cal B},{\bf E}^{2^{t-2}}{\cal B},{\bf E}^{2^{k-1}+2^{t-2}}{\cal A},{\bf E}^{2^{k-1}}{\cal A},{\bf E}^{2^{k-1}+2^{t-2}+1}{\cal A},$ (6)
    ${\bf E}^{2^{k-1}}{\cal A},{\bf E}^{2^{k-1}+2^{t-2}+2}{\cal A},{\bf E}^{2^{k-1}}{\cal A},\ldots,{\bf E}^{2^{k-1}}{\cal A},{\bf E}^{2^{k-1}+2^{t-1}-2}{\cal A},,{\bf E}^{2^{k-1}}{\cal A},{\bf E}^{2^{k-1}+2^{t-1}-1}{\cal A},{\bf E}^{2^{k-1}}{\cal A}]~{}.$
    The same analysis as in Case 1 is done in this case, taken into account that $t=k$. We omit the specific discussions as a more comprehensive analysis will be made in Case 3, which is a generalization of this case. Hence, also in this case we have that $\mathbb{C}$ is a $(2^{k},2^{t};n,2)$-DBAC.
    Case 3: If $12\leq t<n-$ 1, then we have $2^{n-t+1}$ codewords. The first four codewords are defined exactly as in Eqs. (3), (4), (5), and (6).
    Note, that in these four codewords if a pair $({\bf E}^{j}{\cal A},{\bf E}^{j+i}{\cal A})$ was not used in one of these codewords, then also the pair $({\bf E}^{j+i}{\cal A},{\bf E}^{j}{\cal A})$ was not used in these codewords. The same it true for $({\bf E}^{j}{\cal B},{\bf E}^{j+i}{\cal B})$ and $({\bf E}^{j+i}{\cal B},{\bf E}^{j}{\cal B})$. As for the other pairs, if the pair $({\bf E}^{j}{\cal A},{\bf E}^{j+i}{\cal B})$ is contained in these codewords, then the pair $({\bf E}^{j+i}{\cal B},{\bf E}^{j}{\cal A})$ is also contained in these codewords.
    The analysis given for Case 1 also holds in the current case. Adding to it the pairs which are contained in the third and the fourth codewords defined in Eqs. (5) and (6), we have that the pairs which are contained in the four codewords are of the form $({\bf E}^{j}{\cal A},{\bf E}^{j+i}{\cal A})$, $({\bf E}^{j}{\cal B},{\bf E}^{j+i}{\cal B})$, $({\bf E}^{j}{\cal A},{\bf E}^{j+i}{\cal B})$, and $({\bf E}^{j}{\cal B},{\bf E}^{j+i}{\cal A})$, where $-2^{t-1}<i<2^{t-1}$ (in Case 2 we have that $t=k$ and hence also the pair with $i=-2^{t-1}=2^{t-2}$ is contained). Moreover, the pairs $({\bf E}^{j}{\cal A},{\bf E}^{j+2^{k-1}}{\cal A})$ and $({\bf E}^{j}{\cal B},{\bf E}^{j+2^{k-1}}{\cal B})$ are contained in the third codeword. The pairs $({\bf E}^{j}{\cal A},{\bf E}^{j+2^{k-1}}{\cal B})$ and $({\bf E}^{j}{\cal B},{\bf E}^{j+2^{k-1}}{\cal A})$ are contained in the fourth codeword.
    Now, it is straightforward to define the remaining codewords. There are $22^{n-t}-$ 2 codewords of the form
    $[{\cal A},{\bf E}^{i_{1}}{\cal B},{\cal A},{\bf E}^{i_{2}}{\cal B},{\cal A},{\bf E}^{i_{3}}{\cal B},\ldots,{\cal A},{\bf E}^{i_{2^{t-1}}}{\cal B}],$ (7)
    where the $i_{j}$’s in all the $22^{n-t}-$ 2 codewords are all the distinct integers between $2^{t-1}$ and $2^{k}-2^{t-1}$ in any order, except for $2^{k-1}$.
    Finally, first there are $12^{n-t-1}-$ 1 codewords of the form
    $[{\cal A},{\bf E}^{i_{1}}{\cal A},{\cal A},{\bf E}^{i_{2}}{\cal A},{\cal A},{\bf E}^{i_{3}}{\cal A},\ldots,{\cal A},{\bf E}^{i_{2^{t-1}}}{\cal A}]$ (8)
    where the $i_{j}$’s in all the $12^{n-t-1}-$ 1 codewords are all the distinct integers between $2^{t-1}$ and $2^{k}-2^{t-1}$ in any order, except for $2^{k-1}$.
    Secondly, there are also $12^{n-t-1}-$ 1 codewords of the form
    $[{\cal B},{\bf E}^{i_{1}}{\cal B},{\cal B},{\bf E}^{i_{2}}{\cal B},{\cal B},{\bf E}^{i_{3}}{\cal B},\ldots,{\cal B},{\bf E}^{i_{2^{t-1}}}{\cal B}]$ (9)
    where the $i_{j}$’s in all the $12^{n-t-1}-$ 1 codewords are all the distinct integers between $2^{t-1}$ and $2^{k}-2^{t-1}$ in any order, except for $2^{k-1}$.
    Thus, $\mathbb{C}$ is a $(2^{k},2^{t};n,2)$-DBAC.
  2. 2.
    For $k<n-1$, the PF$(n,k)$ contains $2^{n-k}\geq 4$ sequences, ${\cal A}_{1},{\cal A}_{2},\ldots,{\cal A}_{2^{n-k}}$. These sequences are ordered in pairs $\{{\cal A}_{i},{\cal A}_{i+1}\}$, where $i$ is an odd integer between 1 and $12^{n-k}-$ 1. For each pair we construct in $\mathbb{C}$ the same codewords as was done in Eqs. (3)-(9) for ${\cal A}$ and ${\cal B}$. The analysis of the obtained pairs is done exactly as was done in Case 3.
    It remains to form the pairs between ${\cal A}_{i}$ and ${\cal A}_{j}$, where $i<j$ and $\{{\cal A}_{i},{\cal A}_{j}\}\neq\{A_{2\ell+1},A_{2\ell+2}\}$, $10\leq\ell\leq 2^{n-k-1}-$ 1. For each such pair $\{{\cal A}_{i},{\cal A}_{j}\}$ (for simplicity this pair is denoted by $\{{\cal A},{\cal B}\}$), we form the following $2^{k-t+1}$ codewords (when $k\geq t$).
    $[{\cal A},{\bf E}^{i_{1}}{\cal B},{\cal A},{\bf E}^{i_{2}}{\cal B},{\cal A},{\bf E}^{i_{3}}{\cal B},\ldots,{\cal A},{\bf E}^{i_{2^{t-1}}}{\cal B}],$
    where the $i_{j}$’s in all the $2^{k-t+1}$ codewords are all the distinct integers between 00 and $12^{k}-$ 1 in any order.
    This implies that $\mathbb{C}$ is a $(2^{k},2^{t};n,2)$-DBAC.
Thus, for each set of parameters a $(2^{k},2^{t};n,2)$-DBAC, where $k<n<2^{k}$ and $12\leq t\leq 2n-k-$ 1, was constructed. ∎
@@
If $k<n<2^{k}$ and $12\leq t\leq 2n-k-$ 1, then there exists a $(2^{k},2^{t};n,2)$-DBAC.
Theorem 6 does not cover the case where $k=n$, i.e., when the perfect factor contains only a span $n$ de Bruijn sequence. The technique that was used will not work in this case since $\sum_{i=0}^{2^{n}-1}i\equiv 2^{n-1}~{}({\mbox{mod}}~{}2^{n})$ and hence we can have that in at least one of the constructed arrays, the sum of all the shifts will be different from 0 modulo $2^{n}$ as required.
The next theorem and its construction is a generalization of a technique presented by Ma [31], where in [31] one sequence over $\mathbb{Z}_{2^{m}}$ is used, while we use all the sequences of a perfect factor over $\mathbb{Z}_{2^{m}}$.
Theorem 7
[…]
Proof:
@@
The generated codeword form a $2^{n}\times 2^{t}$ matrix and since $\sum_{i=0}^{2^{t}-1}q_{i}\equiv 0~{}({\mbox{mod}}~{}2^{n})$, it follows that the matrix is doubly periodic. The perfect factor ${\cal F}$ has $2^{nm-t}$ sequences of length $2^{t}$ and each $m$-tuple is contained as a window of length $m$ in exactly one of the sequences of ${\cal F}$. From each sequence of ${\cal F}$ we generate one codeword in $\mathbb{C}$ and hence the total number of windows in the codewords of $\mathbb{C}$ is $2^{nm-t}\cdot 2^{n}\cdot 2^{t}=2^{n(m+1)}$. Therefore, to prove that the generated code is a $(2^{n},2^{t};n,m+1)$-DBAC it is sufficient to show that each $n\times(m+1)$ matrix is a window in one of the codewords of $\mathbb{C}$. Let $A$ be such an $n\times(m+1)$ matrix. Each column of $A$ is an $n$-tuple in ${\cal S}$. Hence, we have that $A$ is contained in the $2^{n}\times(m+1)$ vertically periodic array
@@
$\beta=(b_{1},b_{2}-b_{1},b_{3}-b_{2},b_{4}-b_{3},\ldots,b_{m}-b_{m-1})=(p_{j},p_{j+1},\ldots,p_{j+m-1}).$
Since each $m$-tuple is contained in a window of length $m$ in one sequence of ${\cal F}$, it follows that $\beta$ is contained as such a window in such a sequence. The associated codeword in $\mathbb{C}$ has the following $m+1$ successive columns

### v2

@@ III-B Codes from two-dimensional FSRs
The code $\mathbb{C}$ which contains the cycles of the two-dimensional FSR defined in Construction 1 is a $(2^{k},2^{t};n,m)$-DBAC of size $2^{nm-k-t}$.
Proof:
First, observe that the cycles of $\mathbb{C}$ in the graph $G_{{\cal F},m}$ are of length $2^{t}$. (this can be observed by noting that the sum of the indices in a cycle $i_{r}$ is always congruent to $1~{}({\mbox{mod}}~{}2^{n-k})$ and the sum of their shifts is always congruent to $0~{}({\mbox{mod}}~{}2^{k})$ since $t\geq k$).
We start by showing that the horizontal period of each cycle defined in $\mathbb{C}$ is ${m+1=2^{t}}$ and not smaller. If the period of some cycle ${\cal C}$ in $\mathbb{C}$ is smaller than $2^{t}$, i.e., the associated matrix has a smaller horizontal periodicity, then it should be a divisor of $2^{t}$, i.e., a power of 2. Assume to the contrary that such a matrix has a horizontal period $2^{s}$ for some $s$, ${0\leq s<t}$ which implies that it consists of $2^{t-s}$ disjoint $2^{k}\times 2^{s}$ equal vertically periodic matrices. Since $\sum_{r=1}^{2^{t}}i_{r}\equiv 1~{}({\mbox{mod}}~{}2^{n-k})$, it follows that $\sum_{r=1}^{2^{t}}i_{r}=\alpha 2^{n-k}+1$, for some positive integer $\alpha$, and hence $\sum_{r=1}^{2^{s}}i_{r}=\frac{\alpha 2^{n-k}+1}{2^{t-s}}$. But, $\frac{\alpha 2^{n-k}+1}{2^{t-s}}$ is not an integer, a contradiction. Therefore, all the cycles of $\mathbb{C}$ have no smaller horizontal periodicity, i.e., they form $2^{k}\times 2^{t}$ arrays with horizontal period $2^{t}$.
We continue by computing the size of the code $\mathbb{C}$. Each $X_{i_{r}}$, $1\leq r\leq m$, is taken arbitrarily from ${\cal F}$ and hence it can be chosen in $2^{n-k}$ distinct ways. The sequence $X_{i_{m+1}}$ is determined by the equation ${\sum_{r=1}^{m+1}i_{r}\equiv 1~{}({\mbox{mod}}~{}2^{n-k})}$. The first cycle $X_{i_{1}}$ is taken in its zero state, while the next $m-1$ cycles can be taken in any of their $2^{k}$ possible cyclic shifts. The shift of the last cycle $X_{i_{m+1}}$ is determined by the equation $\sum_{r=2}^{m+1}j_{r}\equiv 0~{}({\mbox{mod}}~{}2^{k})$. Hence, each of the middle $m-1$ cycles of ${\cal F}$ can be chosen in $2^{n-k}2^{k}=2^{n}$ distinct ways. The first cycle can be chosen in $2^{n-k}$ distinct ways and the last cycle and its shift are determined by the first $m$ cycles and their shifts. Therefore, there are altogether $2^{n(m-1)}2^{n-k}=2^{nm-k}$ distinct ways to choose all the cycles and their shifts.
We calculate the number of cycles that are counted more than once in this enumeration. Each constructed cycle ${\cal C}$ and its associated codeword can start with any one of its $m+1=2^{t}$ columns. Such a column, say the $r$-th column, will be taken in the shift where its zero state is the first $n$-tuple. Each other column, say the $\ell$-th column is taken in shift $j_{\ell}-j_{r}$, i.e, ${\bf E}^{j_{\ell}-j_{r}}X_{i_{\ell}}$ and hence, the associated matrix is shifted vertically by $i_{r}$ compared to ${\cal C}$ (and horizontally by $r-1$). Therefore, each codeword ${\cal C}\in\mathbb{C}$ is counted exactly $2^{t}$ times in this calculation. Thus, the size of $\mathbb{C}$ is $2^{nm-k}/2^{t}=2^{nm-k-t}$. Note, that the sum of all the shifts, in this new arrangement of the first column, will remain congruent to 0 modulo $2^{k}$ since $t\geq k$ and hence the sum of the new shifts will be changed by $j_{r}2^{t}$ modulo $2^{k}$ which is congruent to 0 modulo $2^{k}$, where the $r$-th column of ${\cal C}$ is the new first column.
Therefore, the number of $n\times m$ windows in all the codewords of $\mathbb{C}$ is $2^{nm-k-t}\cdot 2^{k}\cdot 2^{t}=2^{nm}$. Hence, to complete the proof it is sufficient to show that each $n\times m$ matrix appears as a window in a codeword of $\mathbb{C}$. Consider such an $n\times m$ matrix $A$. The columns of $A$ are $n$-tuples in cycles of ${\cal F}$ and hence these cycles can be arranged in a codeword ${\cal C}\in\mathbb{C}$ with all the necessary shifts as required by $A$. This arrangements can be guaranteed by the definition of the first $m=2^{t}-1$ columns in a codeword of $\mathbb{C}$ to form a window with the matrix $A$ in ${\cal C}$, since there are no constraints on the shifts in the construction for the $m-1$ columns which follows the first.
Thus, $\mathbb{C}$ is a $(2^{k},2^{t};n,m)$-DBAC of size $2^{nm-k-t}$. ∎
[…]
Construction 2
Let ${\cal F}$ be a PF$(n,k)$, where ${\cal F}$ does not contain self-dual sequences and ${\cal S}\in{\cal F}$ if and only if $\bar{{\cal S}}\in{\cal F}$. For each $2<k<n<2^{k}$ and $t\geq 1$ define the following two-dimensional FSR
$f(X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{2^{t}}}X_{i_{2^{t}}})=\bar{X}_{i_{1}},$ (4)
where
$X_{i_{r}}\in{\cal F},~{}1\leq r\leq 2^{t},~{}0\leq j_{r}\leq 2^{k}-1,~{}2\leq r\leq 2^{t},$
and each $X_{i_{r}}$ is a column vector and its cyclic shift ${\bf E}^{j_{r}}$ is relative to its zero state. Let $\mathbb{C}$ be the code defined by the cycles of $f$.
Theorem 5
The code $\mathbb{C}$ which contains the cycles of the two-dimensional FSR defined in Construction 2 is a $(2^{k},2^{t+1};n,2^{t})$-DBAC of size $2^{n2^{t}-k-t-1}$.
Proof:
First, observe that the codewords of $\mathbb{C}$ are cycles of length $2^{t+1}$ in the graph $G_{{\cal F},m}$ associated with $2^{k}\times 2^{t+1}$ matrices.
We start by showing that the horizonal period of each codeword in $\mathbb{C}$ is $2^{t+1}$. If the period of some codeword ${\cal C}$ in $\mathbb{C}$ is smaller than $2^{t+1}$, then it must be a divisor of $2^{t}$, i.e., a power of 2. However, this will imply that the sequence $X_{i_{1}}$ is some shift of the sequence $\bar{X}_{i_{1}}$, i.e., this sequence is self-dual which is impossible by our choice of ${\cal F}$. Therefore, all the codewords of $\mathbb{C}$ have no smaller horizontal periodicity and hence they form $2^{k}\times 2^{t+1}$ doubly-periodic arrays.
We continue by computing the size of the code $\mathbb{C}$. Each $X_{i_{r}}$, $1\leq r\leq 2^{t}$, is taken arbitrarily from ${\cal F}$ and hence it can be chosen in $2^{n-k}$ distinct ways. The first cycle $X_{i_{1}}$ is taken in its zero state, while the next $12^{t}-$ 1 cycles can be taken in any of their $2^{k}$ cyclic shifts. The last $2^{t}$ cycles and their shifts are determined by the first $2^{t}$ cycles, i.e., these are complement cycles with the same shifts. Hence, each of the first $2^{t}$ cycles, except for $X_{i_{1}}$, has $2^{n-k}2^{k}=2^{n}$ possible distinct choices. Therefore, there are $2^{n(2^{t}-1)}\cdot 2^{n-k}=2^{n2^{t}-k}$ distinct choices for the codewords in $\mathbb{C}$ and their shifts.
We must calculate the number of codewords of $\mathbb{C}$ which are counted more than once in this enumeration. Each constructed codeword ${\cal C}$ can start with any of its $2^{t+1}$ columns. This column will be taken in the shift where its zero state is the first $n$-tuple. The other columns are defined by the order of the columns in ${\cal C}$ and their shifts are taken exactly as in ${\cal C}$ relative to the first state of the new first column. Therefore, each codeword ${\cal C}$ is counted exactly $2^{t+1}$ times in this calculation. Thus, the size of $\mathbb{C}$ is $2^{n2^{t}-k}/2^{t+1}=2^{n2^{t}-k-t-1}$.
Therefore, the number of $n\times 2^{t}$ windows in the all the codewords of $\mathbb{C}$ is $2^{n2^{t}-k-t-1}\cdot 2^{k}\cdot 2^{t+1}=2^{n2^{t}}$. Hence, to complete the proof it is sufficient to show that each $n\times 2^{t}$ matrix appears as a window in a codeword of $\mathbb{C}$. Consider such a matrix $A$. The columns of $A$ are $n$-tuples in cycles of ${\cal F}$ and hence these cycles can be rearranged in a codeword ${\cal C}$ to form a window equals to the matrix $A$, since there are no constraints in the construction on the order of $2^{t}$ consecutive columns.
Thus, $\mathbb{C}$ is a $(2^{k},2^{t+1};n,2^{t})$-DBAC of size $2^{n2^{t}-k-t-1}$. ∎
[…]
III-C Other direct constructions
The first direct construction can be used to construct $(2^{k},2^{t};n,1)$-DBACs for all possible parameters. The key idea is to notice that in each such array the columns are sequences of length $2^{k}$ such that each $n$-tuple is contained in exactly one of these columns. This implies that if the size of the code is $\Delta$, then the union of the sequences in all the columns of all the $\Delta$ arrays forms a perfect factor in $G_{n}$ and vice versa. This perfect factor has $2^{n-k}$ cycles of length $2^{k}$, i.e., it is a PF$(n,k)$. Moreover, for this perfect factor, $2^{t}$ can be any divisor of $2^{n-k}$. These observations imply the following construction and its associated corollaries.
Construction 3
Let ${\cal F}$ be an PF$(n,k)$ and let $t$ be any integer such that $1\leq t\leq n-k$. Let $\mathbb{C}$ be the code obtained from $2^{n-k-t}$ ${}~{}2^{k}\times 2^{t}$ arrays, where the columns of each array consists of $2^{t}$ distinct sequences of ${\cal F}$ in any chosen shift.
Corollary 1
@@ III-C Other direct constructions
When $t=0$, a $(2^{k},2^{t};n,1)$-DBAC is a perfect factor and when $t=n-k$, a $(2^{k},2^{t};n,1)$-DBAC is a de Bruijn array.
The last direct construction is a generalization of a technique presented by Ma [32], where in [32] one sequence over $\mathbb{Z}_{2^{n}}$ is used, while we use all the sequences of a perfect factor over $\mathbb{Z}_{2^{n}}$.
Construction 4
@@
Theorem 6
Construction 4 yields a $(2^{n},2^{t};n,m+1)$-DBAC.
Proof:
The generated codeword in $\mathbb{C}$ form a $2^{n}\times 2^{t}$ matrix and, since ${\cal S}$ is a cyclic sequence and $\sum_{i=0}^{2^{t}-1}q_{i}\equiv 0~{}({\mbox{mod}}~{}2^{n})$, it follows that the matrix is doubly-periodic. The reason is that ${\bf E}^{\sum_{i=0}^{2^{t}-1}q_{i}}{\cal S}={\cal S}$, ${\bf E}^{\sum_{i=0}^{2^{t}-1}q_{i}+q_{0}}{\cal S}={\bf E}^{q_{0}}{\cal S}$, ${\bf E}^{\sum_{i=0}^{2^{t}-1}q_{i}+q_{0}+q_{1}}{\cal S}={\bf E}^{q_{0}+q_{1}}{\cal S}$, and so on, which implies that after the first $2^{t}$ columns we continue the columns of the array (codeword) from its beginning. The perfect factor ${\cal F}$ has $2^{nm-t}$ sequences of length $2^{t}$ and each $m$-tuple is contained as a window of length $m$ in exactly one of the sequences of ${\cal F}$. From each sequence of ${\cal F}$ we generate one codeword in $\mathbb{C}$ and hence the total number of windows in the codewords of $\mathbb{C}$ is $2^{nm-t}\cdot 2^{n}\cdot 2^{t}=2^{n(m+1)}$. Therefore, to prove that $\mathbb{C}$ is a $(2^{n},2^{t};n,m+1)$-DBAC it is sufficient to show that each $n\times(m+1)$ matrix is a window in one of the codewords of $\mathbb{C}$. Let $A$ be such an $n\times(m+1)$ matrix. Each column of $A$ is an $n$-tuple in ${\cal S}$. Hence, we have that $A$ is contained in the $2^{n}\times(m+1)$ vertically periodic array
@@
$\beta=(b_{1},b_{2}-b_{1},b_{3}-b_{2},b_{4}-b_{3},\ldots,b_{m}-b_{m-1})=(p_{j},p_{j+1},\ldots,p_{j+m-1})$
and $j+m-1$ is taken modulo $2^{t}$ since the sequence of ${\cal F}$ are cyclic. Since each $m$-tuple is contained in a window of length $m$ in one sequence of ${\cal F}$, it follows that $\beta$ is contained as such a window in such a sequence. The associated codeword in $\mathbb{C}$ has the following $m+1$ successive columns
@@
IV Recursive Constructions for de Bruijn Array Codes
For a recursive construction, we are given a $(2^{k},2^{t};n,m)$-DBAC $\mathbb{C}$ with parameters specified in the construction, and a perfect factor ${\cal F}$ with parameters which depend on the parameters of $\mathbb{C}$. The sequences of ${\cal F}$ are used as indicators associated with the columns of the codewords of $\mathbb{C}$, which we use to decide whether to apply ${\bf D}^{-1}_{0}$ or to apply ${\bf D}^{-1}_{1}$. If the weight of the columns of the codewords in $\mathbb{C}$ is even, then applying ${\bf D}^{-1}_{1}$ on these columns yields codewords of the same vertical length, while if their weight is odd, then the vertical length is doubled. Other distinction between the constructions is the horizontal length of the obtained codewords, it can remain the same or increased. Each construction yield codewords with different sizes. Moreover, the number of constructed codewords in the new code vary between the different constructions. This enables to generate codes with various parameters starting with DBACs obtained by the direct constructions and applying various recursive constructions iteratively. Some analysis for these applications will be discussed in Section VI. All the recursive constructions described here are generalizations and modifications of the two constructions presented by Fan, Fan, Ma, and Siu [17] which are applied on a single array.
Construction 5
Let $\mathbb{C}$ be a $(2^{k},2^{m};n,m)$-DBAC of size $\Delta$, where $k,n,m\geq 2$, whose columns are sequences with even weight. Let $A=[{\cal A}_{1}{\cal A}_{2}~{}\cdots~{}{\cal A}_{2^{m}}]$ be any codeword (a $2^{k}\times 2^{m}$ doubly-periodic array) of $\mathbb{C}$. Let
${\cal S}=[s_{1},s_{2},\ldots,s_{2^{m}}]$
be a span $m$ de Bruijn sequence.
For the code $\mathbb{C}^{\prime}$ define a codeword (array) $B$. Let the $i$-th column of $B$ be $B_{i}={\bf D}^{-1}_{s_{i}}{\cal A}_{i}$, where $1\leq i\leq 2^{m}$. Construct such a codeword for each of the $2^{m}$ distinct cyclic shifts of ${\cal S}$, where for each shift ${\bf E}^{j-1}{\cal S}=[s_{j},s_{j+1},\ldots,s_{2^{m}},s_{1},s_{2},\ldots,s_{j-1}]$ a codeword like $B$ is constructed where ${\bf E}^{j-1}{\cal S}$ replaces ${\cal S}$ in the construction of $B$.